# TEn-CATS: Text-Enriched Audio-Visual Video Parsing with Multi-Scale Category-Aware Temporal Graph

### v1

@@ Abstract
Audio-Visual Video Parsing (AVVP) task aims to identify event categories and their occurrence times in a given video with weakly supervised labels. Existing methods typically fall into two categories: (i) designing enhanced architectures based on attention mechanism for better temporal modeling, and (ii) generating richer pseudo-labels to compensate for the absence of frame-level annotations. However, the first type methods treat noisy segment-level pseudo labels as reliable supervision and the second type methods let indiscriminate attention spread them across all frames, the initial errors are repeatedly amplified during training. To address this issue, we propose a method that combines the Bi-Directional Text Fusion (BiT) module and Category-Aware Temporal Graph (CATS) module. Specifically, we integrate the strengths and complementarity of the two previous research directions. We first perform semantic injection and dynamic calibration on audio and visual modality features through the BiT module, to locate and purify cleaner and richer semantic cues. Then, we leverage the CATS module for semantic propagation and connection to enable precise semantic information dissemination across time. Experimental results demonstrate that our proposed method achieves state-of-the-art (SOTA) performance in multiple key indicators on two benchmark datasets, LLP and UnAV-100.
Index Terms:
Audio-Visual Video Parsing, Weakly Supervised Learning, Audio-Visual Learning, Multimodal Signal Processing.
@@ I Introduction
The Audio Visual Video Parsing (AVVP) task aims to not only identify what and when events occur, but also determine which modality -audio, visual, or both- is responsible for detecting the event [1]. As shown in Fig. 1, a key characteristic distinguishes AVVP from other audio-visual tasks [2, 3, 4, 5, 6, 7] is the inherent temporal asynchrony between modalities: an event may exist in one modality but not in another, or the event may occur at different times across modalities. This modality-specific and temporally misaligned characteristics make fine-grained temporal modeling essential for accurately localizing the time of events occurring and assigning modality tags. In addition, AVVP is often performed under a weakly supervised setting, where only video-level labels are available during training, especially when frame-level labels are infeasible or expensive to obtain. This further causes the challenge of learning temporal information and cross-modal interactions without segment-level supervision. Hence, how to fully explore the semantic relationships between video frames with limited data is a challenge for the AVVP task.
@@
Several approaches tried to address these challenges from two aspects. One focuses on using attention-based encoders that model temporal dependencies across segments [1, 8, 9, 10]. The other one explores generating pseudo-labels, often with the help of pre-trained audio-text or vision-language models, to compensate for missing frame-level annotations and guide cross-modal learning [11, 12, 13, 14, 15]. While both directions have shown effectiveness in addressing different aspects of the AVVP task, they still have some limitations. Firstly, all these approaches tend to treat segments in isolation, either relying on rigid temporal structures or assuming pseudo-labels to be semantically accurate. In practice, pseudo-labels often have noise and may propagate errors when used indiscriminately. What’s more, temporal modeling without semantic awareness fails to capture meaningful inter-segment relationships crucial for consistent event reasoning.
@@
To tackle both issues in a unified way, we propose TEn-CATS, which integrates a Bi-Directional Text Fusion (BiT) module and a Category-Aware Temporal Structure (CATS) module. BiT addresses the pseudo-label noise problem: it first derives text embeddings from pseudo labels via CLAP/CLIP, then performs co-attention with audio-visual features and injects global modality semantics to calibrate the embeddings, yielding segment-level representations with reliable cross-modal alignment. CATS makes up the rigidity of previous temporal designs: it constructs a multi-scale temporal graph in which each segment adaptively selects temporal neighbors according to its predicted event category, capturing event-specific durations and patterns. In summary, BiT ensures cross-modal semantic reliability, while CATS provides event-aware temporal reasoning; they together form a coherent framework that explicitly targets the very limitations—noisy supervision and inflexible temporal modeling—observed in existing AVVP methods.
@@
This work is an extended and enhanced version of our previous conference paper TeMTG [16], where we initially proposed an architecture combining text-enhanced features and multi-hop temporal graphs. However, TeMTG employed a fixed-hop design and lacked category-specific modeling, and only use a simple Multi-Layer Perceptron (MLP) [17] to fuse the audio/visual features with text embeddings. In our new TEn-CATS, it replaces the rigid graph with CATS module, which learns hop preferences and decay factors from the predicted event distribution, giving each segment an event-specific temporal receptive field. And we also design the BiT module to calibrates pseudo-text embeddings for effectively suppressing label noise. In addition, we conduct more ablation and visualization experiments on more datasets to further prove the effectiveness and robustness of our new model.
@@
    To the best of our knowledge, this is the first work to introduce a temporal graph neural network into the AVVP task, where we propose the CATS module that captures segment-specific multi-scale dependencies by adaptively selecting temporal hops based on event semantics.
@@
    The BiT module shifts the conventional design paradigm from fixed text-to-feature alignment to feature-aware text calibration. This enables the pseudo labels to dynamically adapt to modality-specific semantic contexts, thereby mitigating semantic noise.
@@
    Proven by experiments, our framework achieves the state-of-the-art (SOTA) on two benchmark datasets, LLP and UnAV-100 datasets across multiple evaluation metrics.
Figure 1: Illustration of the AVVP task.
Figure 2: Overview of our model. Segment-level audio/visual features are extracted by frozen CLAP/CLIP, and segment-level text embeddings are generated from pseudo labels. We introduce BiT for feature-aware text calibration via co-attention, then aggregate the calibrated features with HAN and enhance them with CATS, a multi-scale temporal graph (residual GAT). A gated fusion and MMIL pooling produce audio (Pa), visual (Pv), and joint (P) predictions.
@@ II-A Audio-Visual Video Parsing
AVVP task aims to jointly detect the occurrence, timing, and type of events using weak video-level annotations. The task was introduced by Tian et al. [1], along with the Look, Listen and Parse (LLP) dataset and a baseline framework that integrates a hybrid attention mechanism (HAN) with a multimodal multiple instance learning based pooling strategy (MMIL Pooling) to extract relevant segment-level cues. Subsequent research efforts mainly focus on two complementary directions. The first is on advancing feature representation backbones. For instance, some researchers use different strategies to enhance intra- and inter-modal semantic consistency [8, 18, 19], while others explored hierarchical or multi-scale temporal structures to better capture event durations across time [10, 20]. The second one is to improve supervision signal by generating pseudo-labels. Some works adopt pre-trained audio-text or vision-language models (e.g., CLAP [21] and CLIP [22]) to infer frame-level labels from video-level tags [23, 12, 13]. Other researchers tried to directly fuse text embeddings generated by the text branch of the pre-trained large language models [14, 15]. These efforts have advanced the AVVP task in various directions. However, these approaches still rely on static temporal structures or treat pseudo-labels as reliable ground-truth without addressing their semantic noise. In contrast, our proposed framework tackles both limitations simultaneously, where we design the BiT module to calibrate noisy pseudo labels through modality-aware conditioning, while the CATS module learns category-dependent multi-scale graphs, thereby offering reliable supervision and flexible temporal modeling under weak labels.
@@ II-B Text-guided Audio-Visual Learning
Recent studies increasingly leverage text as a high-level semantic modality to guide audio-visual learning. Text provides explicit supervision, enabling cross-modal alignment and enhancing representation learning [24]. As examples, texts have been integrated into pretraining frameworks via masked modeling, or tri-modal learning to align audio, visual, and textual modalities [25, 26]. This improves generalization across modalities and supports semantic consistency [27, 28]. Other methods adopt large language models or pre-trained text embeddings to ground audio-visual content in semantic space [29, 30]. These approaches enable spatial-temporal localization and content-aware learning without dense labels. In addition, textual queries have been used for source localization and event understanding, aligning natural language descriptions with audio-visual signals via attention-based fusion [31, 32]. Most existing methods depend on static text embeddings or unidirectional alignment (text to feature), which limits adaptability to ambiguous or evolving events. Moreover, reliance on annotated queries introduces bias and noise in weakly-supervised settings. In contrast, our BiT module produces context-adaptive, noise-robust alignments without relying on static text tokens or additional annotation queries.
@@ II-C Graph-based Audio-Visual Learning
Graph-based models have emerged as a powerful paradigm for audio-visual learning, due to their flexibility in capturing cross-modal relationships, temporal dependencies, and structural semantics [33, 34]. Previous works can be divided into three main categories. The first is to model temporal information using the graph structure [35, 36, 37, 38]. Early methods adopted hand-crafted temporal edge connections for short-term temporal modeling [35, 36]. However, recent works utilize spatial-temporal graph neural networks to capture long-range dependencies [37, 38]. The second is to leverage the cross-modal interaction by constructing multimodal graphs with herterogeneous nodes and edges, thus aligning audio and visual semantics through attention or dynamic edge construction [39, 40, 41]. The third is to build task-oriented graph structure. Custom graph designs have been developed for various audio-visual tasks, such as emotion recognition [42], question answering [43], and event localization [44]. These approaches often incorporate semantic priors, scene attributes, or label co-occurrence patterns into the graph structure, enabling task-specific reasoning. Unlike hand-crafted or task-specific graph designs, our CATS model dynamically selects the jump size based on the predicted event class.
@@ III-A Problem Formulation
In the weakly supervised AVVP task, each input video $\mathcal{S}$ is divided into $T$ segments, denoted as $\mathcal{S}=\{\mathbf{a}_{t},\mathbf{v}_{t}\}_{t=1}^{T}$, where $\mathbf{a}_{t}\in\mathbb{R}^{d}$ and $\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the audio and visual features of the $t$-th segment, with $d$ being the corresponding feature dimensions. The task has two main objectives, one is to identify three types of events: audio-only events $\mathbf{y}_{t}^{a}\in\{0,1\}^{C}$, visual-only events $\mathbf{y}_{t}^{v}\in\{0,1\}^{C}$, and audio-visual events $\mathbf{y}_{t}^{av}\in\{0,1\}^{C}$, where $C$ is the total number of event categories. Each label vector is multi-hot, allowing multiple events to co-occur at the same segment. In addition, audio-visual event is considered to occur only when the audio event and visual event happen simultaneously, which means $\mathbf{y}_{t}^{av}=\mathbf{y}_{t}^{a}\odot\mathbf{y}_{t}^{v}$. The second objective is to localize the temporal boundaries of each event. During training, only weak (i.e. video-level) labels $\mathbf{y}\in\{0,1\}^{C}$ are available.
@@ III-B Overview of The Proposed Framework
We build upon CoLeaF [19] as our baseline and introduce a novel multimodal optimization framework that integrates Bi-Directional text fusion and category-aware temporal structure modeling. Our framework establishes a mutually reinforcing pipeline where semantics-aware feature calibration directly informs adaptive temporal reasoning, enabling a more coherent and context-driven parsing of audio-visual events.  As shown in Fig. 2, we first extract audio and visual features using pre-trained CLAP [21] and CLIP [22] encoders, and generate segment-level text embeddings from pseudo labels provided by VALOR [11]. Then, the Bi-Directional text fusion (BiT) module (Section III-C) is designed to improve the semantic alignment between the modality features and the text features. The fused features are then processed by a HAN [1] with self- and cross-attention. To model event-specific temporal dependencies, we also introduce the CATS module (Section III-D) that builds a learnable graph across segments. A gated fusion module adaptively integrates the outputs from HAN and CATS. Finally, MMIL pooling [1] is applied to produce audio, visual, and joint predictions. Similar to CoLeaF [19], we insert our BiT and CATS modules into both the anchor and reference branches for feature aggregation.
III-C Bi-Directional Text Fusion Module (BiT)
@@
To improve the semantic alignment between the features from audio-visual modalities and the supervision information from weak event labels, we introduce the BiT module that enables fine-grained interaction between audio and visual features and textual guidance at segment level. Before applying the BiT module, we first utilize the segment-level pseudo labels provided by VALOR [11], where each label is a 25-dimensional binary vector indicating the presence ($1$) or absence (0) of event categories. For each segment, we identify the active event classes (i.e., entries with value $1$) and generate the corresponding textual prompts to construct semantic embeddings for both audio and visual modalities. For that, we design different prompt templates. For audio modality, we use the sentence “This is the sound of [event]”; for visual modality, we use the sentence “This is a photo of [event]”. If no event is active for a given segment (i.e., all-zero vector), the prompt “There is no event in this segment” is used for both modalities.
@@
We convert the segment-level pseudo labels into text embeddings using the frozen text encoders from the pre-trained CLAP [21] and CLIP [22]. Then, we use BiT to fuse the modality features $\mathbf{F}_{m}$ and text embeddings $\mathbf{F}_{e}$. As shown in Fig. 3, the bidirectional cross-attention mechanisms are first applied to fuse the modality features and the text embeddings. Each cross-attention output is processed by performing a dropout operation, followed by residual addition and layer normalization to stabilize training. These updates are as follows:
$\mathbf{F}^{{}^{\prime}}_{m}=\text{LN}(\mathbf{F}_{m}+\text{Dropout}(\text{MHA}(\mathbf{F}_{m},\mathbf{F}_{e},\mathbf{F}_{e})))$ (1)
$\mathbf{F}^{{}^{\prime}}_{e}=\text{LN}(\mathbf{F}_{e}+\text{Dropout}(\text{MHA}(\mathbf{F}_{e},\mathbf{F}_{m},\mathbf{F}_{m})))$ (2)
@@
Previous works often rely only on pseudo-label supervision for feature-text alignment, while they are prone to noise and semantic drift. We argue that the modality features contain more reliable semantic cues. As “semantic anchors”, modality features can be better leveraged to refine text representations and mitigate the noise inherent in pseudo-labels. Hence, we utilize a self-attentive pooling mechanism to extract global semantic information from modality features. Self-attentive pooling is performed as follows. Firstly, we calculate an attention score $s$ for each segment through a linear projection. Concretely, the attention weights $\alpha_{t}$ for each segment are computed as:
$\alpha_{t}=\frac{\text{exp}({\mathbf{W}^{\top}_{att}}\mathbf{f}^{{}^{\prime}}_{m}(t))}{{\textstyle\sum_{i=1}^{T}}({\mathbf{W}^{\top}_{att}}\mathbf{f}^{{}^{\prime}}_{m}(i))}\in\mathbb{R}^{b\times 1}$ (3)
where $\mathbf{f}^{{}^{\prime}}_{m}(t)$ means the features in each segment, and $\mathbf{W}_{att}\in\mathbb{R}^{d\times 1}$ is a learnable projection parameter. Afterwards, the global semantic vector $g_{m}$ is computed as the weighted sum of the segment-level features:
$\mathbf{g}_{m}=\sum_{t=1}^{T}\alpha_{t}\cdot\mathbf{f}^{{}^{\prime}}_{m}(t)\in\mathbb{R}^{b\times d}$ (4)
This global representation $g_{m}$ can capture the salient modality-specific semantics, which are subsequently expanded and fused with text embeddings. Finally, the concatenated features are fed into an MLP [17] to project the fused representations back to the original feature with dimension $d$. This enables effective integration of modality-specific semantics and refined textual guidance, producing enhanced segment-level features for downstream event parsing. Note that the BiT module is incorporated into the audio and visual branches separately, while their parameters are shared, which can improve cross-modal training.
Figure 3: Overview of the BiT module. Modality features and text embeddings interact via bi-directional attention, followed by semantic pooling and fusion through an MLP.
III-D Category-Aware Multi-Scale Temporal Graph (CATS)
To model event-specific temporal structures, we propose a category-aware multi-scale temporal graph, which dynamically constructs the temporal graph guided by predicted category information.

### v2

# TEn-CATG: Text-Enriched Audio-Visual Video Parsing with Multi-Scale Category-Aware Temporal Graph

@@ Abstract
Audio-visual video parsing (AVVP) aims to detect event categories and their temporal boundaries in videos, typically under weak supervision. Existing methods mainly focus on (i) improving temporal modeling using attention-based architectures or (ii) generating richer pseudo-labels to address the absence of frame-level annotations. However, attention-based models often overfit noisy pseudo-labels, leading to cumulative training errors, while pseudo-label generation approaches distribute attention uniformly across frames, weakening temporal localization accuracy. To address these challenges, we propose TEn-CATG, a text-enriched AVVP framework that combines semantic calibration with category-aware temporal reasoning. More specifically, we design a bi-directional text fusion (BiT) module by leveraging audio-visual features as semantic anchors to refine text embeddings, which departs from conventional text-to-feature alignment, thereby mitigating noise and enhancing cross-modal consistency. Furthermore, we introduce the category-aware temporal graph (CATG) module to model temporal relationships by selecting multi-scale temporal neighbors and learning category-specific temporal decay factors, enabling effective event-dependent temporal reasoning. Extensive experiments demonstrate that TEn-CATG achieves state-of-the-art results across multiple evaluation metrics on benchmark datasets LLP and UnAV-100, highlighting its robustness and superior ability to capture complex temporal and semantic dependencies in weakly supervised AVVP tasks.
@@ I Introduction
The audio visual video parsing (AVVP) task aims to not only identify what and when events occur, but also determine which modality -audio, visual, or both- is responsible for detecting the event [1]. As shown in Fig. 1, a key characteristic distinguishes AVVP from other audio-visual tasks [2, 3, 4, 5] is the inherent temporal asynchrony between modalities: an event may exist in one modality but not in another, or the event may occur at different times across modalities. This modality-specific and temporally misaligned characteristics make fine-grained temporal modeling essential for accurately localizing the time of events occurring and assigning modality tags. In addition, AVVP is often performed under a weakly supervised setting, where only video-level labels are available during training, especially when frame-level labels are infeasible or expensive to obtain. This further increases the challenge of learning temporal information and cross-modal interactions without segment-level supervision. Hence, how to fully explore the semantic relationships between video frames with limited data is a challenge for the AVVP task.
Figure 1: Illustration of the AVVP task. An event may occur in one modality, or across different modalities, and at different times.
@@
Several approaches have been proposed to address these challenges in two aspects. One focuses on using attention-based encoders that model temporal dependencies across segments [1, 6, 7, 8]. The other explores the generation of pseudo-labels, often with the help of pre-trained audio-text or vision-language models, for the frame missing annotations, and guides cross-modal learning [9, 10, 11, 12, 13]. Although both directions have shown effectiveness in addressing different aspects of the AVVP task, they still have limitations. Firstly, these approaches treat segments in isolation, either relying on rigid temporal structures or assuming pseudo-labels to be semantically accurate. In practice, pseudo-labels are often noisy or inaccurate, and may propagate errors when used indiscriminately. Moreover, temporal modeling without semantic awareness can be limited for capturing meaningful inter-segment relationships crucial for consistent event reasoning.
@@
To address both issues in a unified way, we propose TEn-CATG, a text-enriched (TEn) AVVP framework that integrates a bi-directional text (BiT) fusion module and a category-aware temporal graph (CATG) module. The BiT module is designed to address the problem of pseudo-label noise. More specifically, it first derives text embeddings from pseudo labels via a contrastive language audio pretraining (CLAP) [14] or contrastive language image pretraining (CLIP) [15] model, then performs co-attention with audio-visual features and injects global modality semantics to calibrate the embeddings, yielding segment-level representations with reliable cross-modal alignment. The CATG module addresses the limitation of temporal modelling in previous designs, by constructing a multi-scale temporal graph, where each segment adaptively selects temporal neighbors according to its predicted event category, capturing event-specific durations and patterns. In short, BiT ensures cross-modal semantic reliability, while CATG provides event-aware temporal reasoning. They together form a coherent framework that explicitly addresses the limitations in existing AVVP methods, namely, noisy supervision and inflexible temporal modeling.
@@
This work is an extended version of the text-enhanced multi-hop temporal graph (TeMTG) model [16], presented in our previous conference paper. TeMTG used a fixed-hop design and lacked category-specific modeling, and only a simple multi-layer perceptron (MLP) [17] to fuse audio-visual features with text embeddings. In our new TEn-CATG model, however, the rigid graph is replaced by the CATG module, which learns hop preferences and decay factors from the predicted event distribution, giving each segment an event-specific temporal receptive field. We also design the BiT module to calibrate pseudo-text embeddings for effectively suppressing label noise. In addition, we conducted comprehensive ablation and visualization experiments on additional datasets to demonstrate the effectiveness and robustness of our new model.
@@
    To our knowledge, this is the first work to introduce a temporal graph neural network into the AVVP task, where we propose the CATG module that captures segment-specific multi-scale dependencies by adaptively selecting temporal hops based on event semantics.
@@
    We propose the BiT module, which enables the pseudo labels to dynamically adapt to modality-specific semantic contexts, thereby mitigating semantic noise. This shifts the design paradigm from conventional fixed text-to-feature alignment to feature-aware text calibration.
@@
    We perform comprehensive experiments on two benchmark datasets, LLP and UnAV-100, respectively. The results show that our framework achieves state-of-the-art (SOTA) performance in multiple evaluation metrics.
Figure 2: Overview of our model. Segment-level audio/visual features are extracted by frozen CLAP/CLIP, and segment-level text embeddings are generated from pseudo labels. The BiT module is introduced for feature-aware text calibration via co-attention, then the features are aggregated with HAN and enhanced with CATG, a multi-scale temporal graph (i.e. residual GAT). Finally, a gated fusion and MMIL pooling produce audio (Pa), visual (Pv), and joint (P) predictions.
@@ II-A Audio-Visual Video Parsing
The AVVP task aims to jointly detect the occurrence, timing, and type of events using weak video-level annotations. The task was introduced by Tian et al. [1], along with the look, listen and parse (LLP) dataset and a baseline framework that integrates a hybrid attention mechanism (HAN) with a multimodal multiple instance learning (MMIL) based pooling strategy to extract relevant segment-level cues. Subsequent research efforts focus mainly on two complementary directions. The first is on advancing feature representation backbones, such as the various strategies to enhance intra- and inter-modal semantic consistency [6, 18, 19], and the hierarchical or multi-scale temporal structures to capture event durations across time [8, 20]. The second one is to improve supervision signal by generating pseudo-labels, such as using the pre-trained audio-text or vision-language models (e.g., CLAP [14] and CLIP [15]) to infer frame-level labels from video-level tags [21, 10, 11], and incorporating text embeddings generated by the pre-trained large language models [12, 13]. However, these approaches rely on static temporal structures or treat pseudo-labels as reliable ground-truth without considering their semantic noise. In contrast, we design the BiT module to calibrate noisy pseudo labels through modality-aware conditioning, and the CATS module to learn the category-dependent multi-scale graphs, thereby offering reliable supervision and flexible temporal modeling under weak labels.
@@ II-B Text-guided Audio-Visual Learning
Recent studies increasingly leverage text as a high-level semantic modality to guide audio-visual learning. Text provides explicit supervision, enabling cross-modal alignment and enhancing representation learning [22]. As examples, texts have been integrated into pretraining frameworks via masked modeling, or tri-modal learning to align audio, visual, and textual modalities [23, 24]. This improves generalization across modalities and supports semantic consistency [25, 26]. Other methods adopt large language models or pre-trained text embeddings to ground audio-visual content in semantic space [27, 28]. These approaches enable spatial-temporal localization and content-aware learning without dense labels. In addition, textual queries have been used for source localization and event understanding, aligning natural language descriptions with audio-visual signals via attention-based fusion [29, 30]. These methods rely on static text embeddings or unidirectional alignment (text to feature), which limits adaptability to ambiguous or evolving events. Moreover, reliance on annotated queries introduces bias and noise in weakly-supervised settings. In contrast, our BiT module produces context-adaptive, noise-robust alignments without relying on static text tokens or additional annotation queries.
@@ II-C Graph-based Audio-Visual Learning
Graph-based models have emerged as a powerful paradigm for audio-visual learning, due to their flexibility in capturing cross-modal relationships, temporal dependencies, and structural semantics [31, 32]. Previous works can be divided into three main categories. The first is to model temporal information using the graph structure [33, 34, 35, 36], employing hand-crafted temporal edge connections for short-term temporal modeling [33, 34], or using spatial-temporal graph neural networks to capture long-range dependencies [35, 36]. The second is to leverage the cross-modal interaction by constructing multimodal graphs with herterogeneous nodes and edges, thus aligning audio and visual semantics through attention or dynamic edge construction [37, 38]. The third is to customize a graph structure for a variety of audio-visual tasks such as emotion recognition [39], question answering [40], and event localization [41]. These approaches often incorporate semantic priors, scene attributes, or label co-occurrence patterns into the graph structure, enabling task-specific reasoning. Unlike hand-crafted or task-specific graph designs, our CATG model selects adaptively the jump size based on the predicted event class.
@@ III-A Problem Formulation
In the weakly supervised AVVP task, each input video $\mathcal{S}$ is divided into $T$ segments, denoted as $\mathcal{S}=\{\mathbf{a}_{t},\mathbf{v}_{t}\}_{t=1}^{T}$, where $\mathbf{a}_{t}\in\mathbb{R}^{d}$ and $\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the audio and visual features of the $t$-th segment, with $d$ being the corresponding feature dimensions. The task has two main objectives, one is to identify three types of events: audio-only events $\mathbf{y}_{t}^{a}\in\{0,1\}^{C}$, visual-only events $\mathbf{y}_{t}^{v}\in\{0,1\}^{C}$, and audio-visual events $\mathbf{y}_{t}^{av}\in\{0,1\}^{C}$, where $C$ is the total number of event categories. Each label vector is multi-hot, allowing multiple events to co-occur at the same segment. In addition, audio-visual event is considered to occur only when the audio event and visual event happen simultaneously, which means $\mathbf{y}_{t}^{av}=\mathbf{y}_{t}^{a}\odot\mathbf{y}_{t}^{v}$, where $\odot$ represents the logical and operation. The second objective is to localize the temporal boundaries of each event. During training, only weak (i.e. video-level) labels $\mathbf{y}\in\{0,1\}^{C}$ are available.
@@ III-B Overview of The Proposed Framework
We build upon CoLeaF [19] as our baseline and introduce a novel multimodal optimization framework that integrates bi-directional text fusion and category-aware temporal structure modeling. Our framework establishes a mutually reinforcing pipeline where semantics-aware feature calibration directly informs adaptive temporal reasoning, enabling a more coherent and context-driven parsing of audio-visual events. As shown in Fig. 2, we first extract audio and visual features using pre-trained CLAP [14] and CLIP [15] encoders, and generate segment-level text embeddings from pseudo labels provided by VALOR [9]. Then, the bi-directional text fusion (BiT) module (Section III-C) is designed to improve the semantic alignment between the modality features and the text features. The fused features are then processed by a HAN [1] with self- and cross-attention. To model event-specific temporal dependencies, we also introduce the CATG module (Section III-D) that builds a learnable graph across segments. A gated fusion module adaptively integrates the outputs from HAN and CATG. Finally, MMIL pooling [1] is applied to produce audio, visual, and joint predictions. Similar to CoLeaF [19], we insert our BiT and CATG modules into both the anchor and reference branches for feature aggregation.
III-C Bi-Directional Text Fusion
@@
To improve the semantic alignment between the features of audio-visual modalities and the supervision information from weak event labels, we introduce the BiT module that enables fine-grained interaction between audio and visual features and textual guidance at segment level. Before applying the BiT module, we first utilize the segment-level pseudo labels provided by VALOR [9], where each label is a 25-dimensional binary vector indicating the presence ($1$) or absence (0) of event categories. For each segment, we identify the active event classes (i.e., entries with value $1$) and generate the corresponding textual prompts to construct semantic embeddings for both audio and visual modalities. For that, we design different prompt templates. For audio modality, we use the sentence “This is the sound of [event]”; for visual modality, we use the sentence “This is a photo of [event]”. If no event is active for a given segment (i.e., all-zero vector), the prompt “There is no event in this segment” is used for both modalities.
@@
We convert the segment-level pseudo labels into text embeddings using the frozen text encoders from the pre-trained CLAP [14] and CLIP [15]. Then, we use BiT to fuse the modality features $\mathbf{M}\in\mathbb{R}^{b\times T\times d}$ and text embeddings $\mathbf{E}\in\mathbb{R}^{b\times T\times d}$, where $b$ represents batch size. As shown in Fig. 3, the bidirectional cross-attention mechanisms are first applied to fuse the modality features and the text embeddings. Each cross-attention output is processed by performing a dropout operation, followed by residual addition and layer normalization to stabilize training. These updates are as follows:
$\mathbf{M}^{{}^{\prime}}=\text{LN}(\mathbf{M}+\text{Dropout}(\text{MHA}(\mathbf{M},\mathbf{E},\mathbf{E})))$ (1)
$\mathbf{M}^{{}^{\prime}}=\text{LN}(\mathbf{M}+\text{Dropout}(\text{MHA}(\mathbf{E},\mathbf{M},\mathbf{M})))$ (2)
@@
Previous works, e.g.  [21, 10, 11], often rely only on the supervision of pseudo-labels for feature-text alignment. However, they are prone to noise and semantic drift. We argue that the modality features contain more reliable semantic cues. As “semantic anchors”, modality features can be better leveraged to refine text representations and mitigate the noise inherent in pseudo-labels. Hence, we utilize a self-attentive pooling mechanism to extract global semantic information from modality features. Self-attentive pooling is performed as follows. Firstly, we calculate an attention score $s$ for each segment through a linear projection. Concretely, the attention weights $\boldsymbol{\alpha}_{t}$ for each segment are computed as:
$\boldsymbol{\alpha}_{t}=\frac{\text{exp}({\mathbf{w}^{\top}_{att}}\mathbf{M}^{{}^{\prime}}_{t})}{{\textstyle\sum_{i=1}^{T}}({\mathbf{w}^{\top}_{att}}\mathbf{M}^{{}^{\prime}}_{i})}\in\mathbb{R}^{B}$ (3)
where $\mathbf{M}^{{}^{\prime}}_{t}$ means the features corresponding to each segment $t$ in $\mathbf{M}^{{}^{\prime}}$, and $\mathbf{w}_{att}\in\mathbb{R}^{d}$ is a learnable projection parameter. Afterwards, the global semantic vector $\mathbf{g}_{m}$ is computed as the weighted sum of the segment-level features:
$\mathbf{G}=\sum_{t=1}^{T}\alpha_{t}\cdot\mathbf{M}^{{}^{\prime}}_{t}\in\mathbb{R}^{B\times}$ (4)
This global representation $\mathbf{G}$ can capture the salient modality-specific semantics, which are subsequently expanded and fused with text embeddings. Finally, the concatenated features are fed into an MLP [17] to project the fused representations back to the original feature with dimension $d$. This enables effective integration of modality-specific semantics and refined textual guidance, producing enhanced segment-level features for downstream event parsing. Note that the BiT module is incorporated into the audio and visual branches separately, while their parameters are shared, which can improve cross-modal training.
Figure 3: Overview of the BiT module. The modality features and text embeddings interact via bi-directional attention, followed by semantic pooling and fusion through an MLP.
III-D Category-Aware Multi-Scale Temporal Graph
To model event-specific temporal structures, we propose the CATG module, which dynamically constructs the temporal graph guided by predicted category information.